OpenStudy (anonymous):

medal and fan, please help me ): You originally draw a design for an art contest on a 3 in. x 5 in. card. The second phase of the contest requires the drawing to be transferred to an 8.5 in x 11 in. standard sheet of paper and utilize as much of the space on the paper as possible. You determine that the largest size one of the dimensions of your drawing can be is 10.5 in. What is the length of the other dimension if the two drawings are similar? Type your exact answer in the blank without the units, and round to the nearest tenths.

OpenStudy (anonymous):

@JuanitaM

OpenStudy (anonymous):

I would say 5 * 2.1 = 10.5 so the other side is 3 * 2.1 = 6.3

OpenStudy (anonymous):

alright (: but could you explain?

OpenStudy (anonymous):

Logically, they said that the Max length of the new dimension is 10.5 so b/w 3 and 5 ; the 10.5 MUST relate to 5 10.5 / 5 = 2.1 so old algebra rule - what you do to one, you must do to the other so 2.1 * 3 is how I get the other dimension does that make sense

OpenStudy (anonymous):

And because it has to be rounded, would it be 11 inches and 6 inches?

OpenStudy (anonymous):

oh wait, i'm sorry the number we would be rounding is 2.1 ----> 2 inches?

OpenStudy (anonymous):

No, read the problem carefully, you have to take the 3 by 5 card and use AS MUCH space as possible to get on a 8.5 by 11 page

OpenStudy (anonymous):

round to the nearest tenth

OpenStudy (anonymous):

What is the length of the other dimension if the two drawings are similar? that is the question - that is 2.1 * 3 = 6.3

OpenStudy (anonymous):

So in the reply bar of my assignment I put 2.1 * 3 = 6 ?

OpenStudy (anonymous):